OpenStudy (anonymous):

Solve by completing the square: 9x^2+30x+0=0 and 4x^2-6x=0

OpenStudy (anonymous):

FOR 9x^2+30x+0=0: - divide the middle term 30x by (2*3x) and then square the result. you should get 25 - Add 25 to both sides (3x + 5)^2 = 25 3x + 5 = +/- 5 Then solve for x

OpenStudy (anonymous):

Use same technique for the other one

OpenStudy (anonymous):

Where did (2*3x) come from and how did it get 25?

OpenStudy (anonymous):

so the formula for (a + b)^2 = a^2 + 2ab + b^2 when you use completing the square, you are finding b^2 and what you have is a^2 + 2ab so to find b, u need to divide 2ab by 2a, and then you square the result to get b^2 to find "a", you just take the principal square root of a^2 (first term) IN THIS CASE 30x/2*3x = 30x/6x = 5 = b 5^2 = 25 = b^2

OpenStudy (anonymous):

square root of 9x^2 = 3x